Define the following theorems with respect to a random vector

i) Central limit theorem

ii) Weak law of large numbers


1
Expert's answer
2022-05-11T12:11:40-0400

i). Central limit theorem: Let "X_1,X_2,X_3,\ldots,X_n" be independent and identically distributed random variables with expected value "E[X_i]=\mu<\infty" and variance "Var[X_i]=\sigma^2<\infty". Denote "\bar{X}=\frac{X_1+X_2+\ldots+X_n}{n}", "Z_n=\frac{\bar{X}-\mu}{\frac{\sigma}{\sqrt{n}}}=\frac{X_1+X_2+\ldots+X_n-n\mu}{\sqrt{n}\sigma}". Then, "\lim_{n\rightarrow\infty}P(Z_n\leq x)=F(x)", where "F(x)" is the cumulative distribution function of the standard normal distribution. I.e., "F(x)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^xe^{-\frac{t^2}{2}}dt."

ii). Weak law of large numbers: Let "X_1,X_2,X_3,\ldots,X_n" be independent and identically distributed random variables with expected value "E[X_i]=\mu<\infty" and variance "Var[X_i]=\sigma^2<\infty". Denote "\bar{X}=\frac{X_1+X_2+\ldots+X_n}{n}" Then, for any "\varepsilon>0", "\lim_{n\rightarrow\infty}P(|\bar{X}-\mu|\geq\varepsilon)=0."